day shift, 30 Sept 99

Marj Corcoran (corcoran@physics.rice.edu)
Thu, 30 Sep 1999 15:48:42 -0500 (CDT)


Shift Summary for day shift, 30 Sept 99

On Shift: M. Corcoran and Hogan

Hours of beam available at all intensities: 0
Hours of beam available at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Accelelerator down for entire shift.

========================================================================

Runs ended this shift:
Run 14776 to 14778--cosmic testing, not written to tape

=======================================================================

Tapes brought to FCC this shift: none

==================================================================

Accesses during shift: BobH and Hogan, to examine the loading dock.
BobT and Rick C make access to NM2 to put lead washers around the
90 degree accidental trigger. Marj and DaveN to put alcohol in the
flammables cabinet.

=================================================================
Description of shift: Beam off a 7AM for scheduled down time.
Hogan, Marj, BobT, Jason L. debug problems with spill gate and
funny dip in DC3X illumination for cosmics. The rest of the shift spent trying
to understand the funny dip in DC3X illumination with cosmics.

==================================================================
More Details:

Hogan and Marj discover that the spill gate in rack 160 was not toggling
properly. The problem was the "fake spill cable" installed last night...it
worked OK when there was no spill, but seemed to hang up the spill gate
when there was a real spill signal. Removing the cable installed last night
fixed the problem and the laser CsI monitoring was restored.

Spent a good part of the shift trying to understand the cause for the
"notch" in the DC3X cosmic illumination. We discover that it has something
to do with trigger 16...with Trig16 turned off, the notch goes away. If
Trig16 is on but prescaled by 100, the notch goes away. If Trig16 is on, but
prescaled by 10, the effect is there but smaller. We are baffled.

BobT and Rick C make access to work on 90 degree accidental counter. They
put lead washers around the scintillators to try and reduce the rate.

Peter S does L1 timing scan.

Arun fixes the destination plane logic to include trig16. This should stop
Elliott's complaining about an "outbox problem", but does not fix
the problem in DC3X illumination.

NM2SR reversing switch is fixed. We should wait at least 1.5 minutes
between commands to the reversing switch, otherwise it could lock up.
This is true of all reversing switches.

===================================================================

Evening shift: Please label some UPA, UPB, and UPC tapes. We are
VERY LOW!